OpenStudy (anonymous):

A travel mug is shaped like a cylinder. It is 9 CM wide and 15 CM tall. Find its volume to the nearist tenth

OpenStudy (jdoe0001):

 \(\Large \bf \textit{volume of a cylinder}=\pi r^2 h \\ \quad \\ r = radius=\frac{diameter}{2}\)